Starting from rest, a car accelerates at 2.5m/s2 up a hill that is inclined 5.1 ? above the horizontal.
a) How far horizontally has the car traveled in 14s ?
b) How far vertically has the car traveled in 14s ?
Answer #1

ax = 2.5 cos5.1 = 2.49 m/s^2

ay = 2.5 sin5.1 = 0.22 m/s^2

sx = 1/2 * axt^2

= 1/2 * 2.49*14^2 = 244.02 m

sy = 1/2 * 0.22*14^2 = 21.56 m(answer a)

s = sqrt(244.02^2 + 21.56^2) = 244.97 m (answer a)
Answer #2

a) vertically

s=1/2at^2 sin thetha

s=sin(5.1)*.5*2.5*14^2

=226.8m
b) Hrizontally

s=1/2at^2 cos thetha

Cos[5.1]*.5*2.5*14^2

=92.6 horizontal
Answer #3

here,

a = 2.5 m/s^2

t = 14 s

using second equation of motion ,

d= ut + 0.5 * at^2

d = 0*14 + 0.5 * 2.5 *14^2

d = 254 m

a)
Horizontal distance = 254 * cos(5.1)

Horizontal distance =244 m

car has travelled 244 m horizontally .

b )

vertical distance = 254*sin(5.1)

vertical distance = 22.6 m

it will travel 22.6 m vertically